OpenStudy (anonymous):

m/8 + 4 = 16 what is m?

OpenStudy (anonymous):

96

OpenStudy (anonymous):

96

OpenStudy (anonymous):

how is it 96?

OpenStudy (anonymous):

m/8= 16-4 =12

OpenStudy (katrinakaif):

m/8 = 16 - 4 m/8 = 12 m = 12 * 8 m = 96

OpenStudy (anonymous):

i forgot to times it :P hehe thank you